OpenStudy (anonymous):

What is the product written in scientific notation? (2.1 × 10^–2)(3.2 × 10^–4)

OpenStudy (anonymous):

A. 6.72 × 10^8 B. 6.72 × 10^–6 C. 5.3 × 1^02 D.5.3 × 10^–6 @undeadknight26

undeadknight26 (undeadknight26):

I THINK This would be B...

OpenStudy (anonymous):

u sure o.o cuz this is a SEMESTER assement!

undeadknight26 (undeadknight26):

Yup.

OpenStudy (anonymous):

SOLVE. 

undeadknight26 (undeadknight26):

ummm: 3p/5 = 12 x5 x5 3p = 60 /3 /3 p = 20